OpenStudy (jkbo):

Please help What are all the real and complex roots of the polynomial x^3 − 5x^2 − x + 5, given that one root is x = -1

OpenStudy (anonymous):

It look alike they want you to use synthetic division based on the directions. Do you know how to do that? Though, you don't have to use synthetic since the expression factors. Do you know how to do that? either way will work.

OpenStudy (jkbo):

No I dont know how to do that @gryphon

OpenStudy (anonymous):

To factor you factor the GCF x^2 out of the first two terms, and -1 out the second two terms to get\[x^2(x-5)-(x-5)=0\]Then factor out the x-5 expression\[(x-5)(x^2-1)=0\]set each factor equal to 0:\[x-5=0cupx^2-1=0\]and solve\[x=5, x=1, x=-1\]These are the roots

OpenStudy (anonymous):

Synthetic division approach is a little more work; I'll show you if you want to see it.

OpenStudy (anonymous):

The synthetic division approach